OpenStudy (anonymous):

What is an example of a trigonometric ratio that can be used to prove that a triangle is a right triangle?

OpenStudy (anonymous):

I cannot use the pythagorean theorem for this, and I can't think of any trigonometric ratios to use. Here's the triangle in question: http://i.imgur.com/83dWTH9.png

OpenStudy (triciaal):

Maybe something like Sin^2 + cos ^2 = 1 or ?

OpenStudy (triciaal):

to solve this triangle (3 lengths given) we use the cosine rule which was from the pythagorean theorem the longest side is opposite the largest angle and after that we use the Sin rule to find the other angles Angles is a triangle add up to 180

OpenStudy (triciaal):

Cosine rule c^2 = a^2 + b^2 -2ab Cos C Sin rule a/Sin A = b/Sin B = c/Sin C